# Horizon Brightened Acceleration Radiation from Massive Vector Fields

### v1

@@ IV.3 Counter-rotating channel and selection rules
    Initial $|g;0_{B}\rangle$ forbids the rotating $\propto a_{\nu\ell m\lambda}\sigma_{eg}$ channel; only the counter-rotating $\propto a^{\dagger}_{\nu\ell m\lambda}\sigma_{eg}$ contributes to excitation. This singles out the terms displayed in (57)-(58) and underlies the emission-before-absorption picture adopted in the master-equation treatment.
@@
    $\nu\geq m_{V},\qquad k_{\infty}(\nu)=\sqrt{\nu^{2}-m_{V}^{2}},$ (59)
    so modes below threshold do not contribute to the escaping flux. In the stationary-phase evaluation, the rapidly varying piece depends only on $\nu>0$, thus preserving the same detailed-balance structure as in Eqs. (B13)-(B15) of Scully:2017utk.
@@
    $\sum_{\lambda}\!\big|u\!\cdot\!\varepsilon^{(\lambda)}\big|^{2}=-1+\frac{\Omega^{2}(\tau,\nu)}{m_{V}^{2}},$ (60)
@@
    Since all master fields reduce to plane waves $\sim e^{-i\nu(t-r_{*})}$ as $r\to r_{g}^{+}$, the same stationary-phase structure that yielded Eq. (B14) in Scully:2017utk in the scalar case recurs here, while the mass and polarization affect only the prefactor (and the $\nu\geq m_{V}$ threshold).
These rules ensure that, after summing over $(m,\lambda)$ and implementing the single-mode cavity filter, the excitation probabilities reduce to the forms already prepared in (44)-(45) and (55)-(56). When we convert these amplitudes into rates and feed them into the master equation, we will recover the same detailed-balance ratio $\Gamma_{e}/\Gamma_{a}=e^{-2\xi}$ with $\xi=2\pi r_{g}\nu/c$, exactly mirroring the scalar result while carrying the Proca-specific prefactors and the hard mass threshold.
With the emission amplitudes written using $\nu\left[t(\tau)-r_{*}(\tau)\right]$ in the rapidly varying phase, the stationary-phase evaluation repeats Scully et al.’s steps (their (B11)–(B14)) verbatim Scully:2017utk, yielding the same Planckian factor, while the Proca mass/polarization affects only the smooth prefactor and the $\nu\geq m_{V}$ threshold.
@@ V Excitation probability along the infall worldline
We evaluate the first-order excitation amplitudes derived in Section IV along the infalling geodesic. For both detector models, the integrals are dominated by the near-horizon segment where the phase is governed by the combination $t-r_{*}$. Using the geodesic relations from Section II and the unit-flux mode normalization from Section III, the worldline integral reduces, in the large-$\omega$ regime, to an endpoint (near-horizon) asymptotics that yields a gamma-function. Its modulus square produces the Planckian detailed-balance factor $\left(e^{4\pi\nu}-1\right)^{-1}$. Mass and polarization enter only through smooth prefactors and the asymptotic kinematic threshold $\nu\geq m_{V}$.
@@ V.1 General stationary-phase integral
For a single outgoing mode $(\nu,\ell,m,\lambda)$, the first-order counter-rotating amplitude has the structure
$\mathcal{A}_{\bullet}(\nu,\ell,m,\lambda)=\int_{-\infty}^{+\infty}\!d\tau\;\mathcal{G}_{\bullet}(\tau,\nu,\ell,m,\lambda)\,e^{\,i\Phi(\tau,\nu)},$ (61)
where $\bullet\in{\text{mono},\text{dip}}$ denotes the detector model. The smooth envelope $\mathcal{G}_{\bullet}$ collects the $1/[r(\tau)\sqrt{k_{\infty}(\nu)}]$ factor from mode normalization and the appropriate polarization weight (cf. (45) for the current coupling and its dipole analogue), while the phase is
$\Phi(\tau,\nu)=\omega\tau+\nu\,\left[t(\tau)-r_{*}(\tau)\right]+\text{(constant)}$ (62)
@@
$\frac{dr}{d\tau}=-\frac{1}{\sqrt{r}},\qquad\frac{dt}{d\tau}=\frac{r}{r-1},\qquad\frac{dr_{*}}{d\tau}=\frac{dr}{d\tau}\,\frac{dr_{*}}{dr}=-\frac{\sqrt{r}}{\,r-1\,}.$ (63)
@@
$\frac{d}{d\tau}(t-r_{*})=\frac{r+\sqrt{r}}{r-1}\approx\frac{1+(1+\delta/2)}{\delta}\approx\frac{2}{y}.$ (64)
@@
$t(\tau)-r_{*}(\tau)=-\,2\ln y+C_{h}+\mathcal{O}(y),\qquad y\to 0^{+},$ (65)
consistent with surface gravity $\kappa=\frac{1}{2}$ in our dimensionless units. Substituting (65) into (62) and absorbing constants into an overall phase,
$\Phi(\tau,\nu)=\Phi_{h}-\omega y-2\nu\ln y+\Delta(\nu)\,r_{*}(\tau_{h})+\mathcal{O}(y).$ (66)
Since $\mathcal{G}_{\bullet}(\tau,\nu,\ell,m,\lambda)$ is smooth, we replace it by its horizon value $\mathcal{G}_{\bullet}(\tau_{h};\nu,\ell,m,\lambda)\equiv\mathcal{G}_{\bullet h}(\nu,\ell,m,\lambda)$ at leading order. The amplitude (61) becomes
$\mathcal{A}_{\bullet}(\nu,\ell,m,\lambda)=\mathcal{N}_{\bullet}(\nu,\ell,m,\lambda)\,\underbrace{\int_{0}^{\infty}\!dy\;y^{-2i\nu}\,e^{-i\omega y}}_{\displaystyle\mathcal{I}(\nu,\omega)},$ (67)
@@
The elementary Laplace-type integral in (67) is expressed in terms of the gamma function,
$I(\nu,\omega)=(i\omega)^{-(1-2i\nu)}\Gamma(1-2i\nu),\qquad\omega>0.$ (68)
@@
$|\mathcal{I}(\nu,\omega)|^{2}=\frac{1}{\omega^{2}}\left|\Gamma(1-2i\nu)\right|^{2}=\frac{2\pi\nu}{\omega^{2}}\;\frac{1}{\sinh(2\pi\nu)}=\frac{4\pi\nu}{\omega^{2}}\;\frac{1}{e^{2\pi\nu}-e^{-2\pi\nu}}\;.$ (69)
Equation (69) governs the thermal character of the problem. While $|\mathcal{I}|^{2}$ itself follows a statistics-dependent profile (here proportional to $\text{cosech}(2\pi\nu)$), the crucial feature for thermodynamics is the relation between excitation (emission) and de-excitation (absorption). All dependence on the detector model (current vs. dipole), the Proca mass and polarizations, and the asymptotic threshold $\nu\geq m_{V}$ resides in the smooth prefactor $\mathcal{N}_{\bullet}$ and in $k_{\infty}(\nu)$. Consequently, at leading order in $1/\omega$, the ratio of probabilities satisfies the detailed balance condition $P_{\text{exc}}/P_{\text{abs}}=e^{-4\pi\nu}$ (derived below via the KMS condition), ensuring the steady state is thermal.
@@ V.2 Universality of the thermal factor
The Planckian denominator obtained in (69) is independent of the detector model (current vs. dipole), the field’s spin, the Proca mass, and greybody details. The reason is structural: after isolating a single outgoing Schwarzschild mode, every first-order counter-rotating amplitude reduces to an integral of the form
$\int d\tau\,\underbrace{\mathcal{G}(\tau,\nu)}_{\text{smooth}}\,\exp\!\big\{i\omega\tau+i\nu\,\left[t(\tau)-r_{*}(\tau)\right]+i\Delta(\nu)r_{*}(\tau)\Big\},$ (70)
@@
$|\mathcal{I}(\nu,\omega)|^{2}=\frac{4\pi\nu}{\omega^{2}}\,\frac{1}{e^{4\pi\nu}-1}\;.$ (73)
Equation (73) is the universal thermal factor for the excitation probabilities. All dependence on the detector realization (current vs. dipole), polarization content, Proca mass and threshold $\nu\geq m_{V}$, and barrier transmission enters only through the smooth prefactor that multiplies (73).
The universality of the excitation mechanism is displayed in Fig. 2, which plots the squared modulus of the near-horizon integral $|\mathcal{I}(\nu)|^{2}$. This profile, governed by the $\text{cosech}(2\pi\nu)$ dependence derived in Eq. (69), dictates the thermal response of the detector. Crucially, this curve is identical for both the monopole and dipole couplings and for any mass $m_{V}$, confirming that the thermal character arises solely from the geometric analyticity of the Rindler time coordinate $t-r_{*}$ rather than the specific details of the probe or the field. The detailed balance ratio required for thermal equilibrium emerges from the symmetry of this kernel under $\nu\to-\nu$.
@@
The same near-horizon analyticity that leads to (73) also fixes detailed balance between the counter-rotating excitation channel and its absorption counterpart (appearing later in the master equation). Under $\nu\to-\nu$ the logarithmic phase in (71) implements the (Kubo-Martin-Schwinger) KMS-type periodicity of Rindler time, giving
$P_{\rm abs}(\nu)=e^{\,4\pi\nu}\,P_{\rm exc}(\nu)\;,$ (74)
@@ V.3 Prefactors, thresholds, and detailed balance
The near-horizon integral in (67)-(69) fixes the entire $\omega$- and $\nu$-dependence that carries the thermal factor. Everything else, from detector realization, polarization structure, greybody transmission, and the Proca mass, enters multiplicatively through a smooth prefactor evaluated on the horizon segment, together with the asymptotic kinematic threshold $\nu\geq m_{V}$.
@@
$P_{\text{exc}}^{\bullet}(\nu,\ell)=\frac{4\pi\nu}{\omega^{2}}\;\frac{\mathcal{K}_{\bullet}\,\Xi_{\ell}(\nu)}{e^{4\pi\nu}-1}\;\frac{\Theta(\nu-m_{V})}{k_{\infty}(\nu)}\;\overline{\mathcal{P}}_{\bullet}(\nu,\ell)\left[1+\mathcal{O}\!\left(\tfrac{\nu}{\omega}\right)\right],\qquad k_{\infty}(\nu)=\sqrt{\nu^{2}-m_{V}^{2}},$ (75)
where $\mathcal{K}_{\bullet}$ collects the detector coupling and matrix element $(\mathcal{K}_{\text{mono}}=g_{V}^{2}|m_{eg}|^{2}$, $\mathcal{K}_{\text{dip}}=g_{d}^{2}|d_{eg}|^{2})$, $\Xi_{\ell}(\nu)$ is the single-channel transmission (greybody) through the exterior potential barrier, $k_{\infty}(\nu)^{-1}$ comes from the unit-flux mode normalization at infinity, and $\overline{\mathcal{P}}_{\bullet}(\nu,\ell)$ is a smooth, order-unity polarization/geometry factor fixed by the corresponding contraction in Section IV and by the cavity’s angular selection.
Specializing (75) to the two detector models requires handling the sum over $m$. For an atom falling radially along the symmetry axis ($\theta\approx 0$), the interaction is weighted by $Y_{\ell m}(\theta\approx 0)$, which vanishes for $m\neq 0$. Consequently, the sum over $m$ collapses to the single $m=0$ term. With this selection, the probabilities become
$P_{\text{exc}}^{\text{(mono)}}(\nu,\ell)=\frac{4\pi\nu}{\omega^{2}}\;\frac{g_{V}^{2}|m_{eg}|^{2}}{e^{4\pi\nu}-1}\;\frac{\Xi_{\ell}(\nu)}{k_{\infty}(\nu)}\;\Theta(\nu-m_{V})\;\overline{\mathcal{P}}_{\text{mono}}(\nu,\ell,m=0)\;,$ (76)
$P_{\text{exc}}^{\text{(dip)}}(\nu,\ell)=\frac{4\pi\nu}{\omega^{2}}\;\frac{g_{d}^{2}|d_{eg}|^{2}}{e^{4\pi\nu}-1}\;\frac{\Xi_{\ell}(\nu)}{k_{\infty}(\nu)}\;\Theta(\nu-m_{V})\;\overline{\mathcal{P}}_{\text{dip}}(\nu,\ell,m=0)\;.$ (77)
@@
The hard mass threshold is enforced by $\Theta(\nu-m_{V})$; close to threshold the asymptotic momentum $k_{\infty}(\nu)$ suppresses the probability as $k_{\infty}^{-1}$, while the greybody factor $\Xi_{\ell}(\nu)$ further damps low $\nu$. In the massless limit $m_{V}\to 0$ one has $k_{\infty}\to\nu$ so the prefactor tends to the familiar $1/\nu$ normalization, reproducing the scalar-like spectral scaling at fixed $\Xi_{\ell}(\nu)$ and $\overline{\mathcal{P}}_{\bullet}$.
Detailed balance follows from the same near-horizon analyticity that produced (73). Replacing $\nu\to-\nu$ in the phase reverses the sign of the logarithm and yields
$P_{\text{abs}}^{\bullet}(\nu,\ell)=e^{\,4\pi\nu}\,P_{\text{exc}}^{\bullet}(\nu,\ell),\qquad\nu>0,$ (78)
@@
$\frac{P_{\text{exc}}^{\bullet}(\nu,\ell)}{P_{\text{abs}}^{\bullet}(\nu,\ell)}=\frac{1}{e^{\,4\pi\nu}}\;.$ (79)
Equations (76)-(79) are the inputs we will promote to rates in the next section by dividing by the interaction time set by the cavity transit and by incorporating the single-mode selection, thereby closing the loop to the master-equation description.

### v2

@@ IV.3 Counterrotating channel and selection rules
    Initial $|g;0_{B}\rangle$ forbids the rotating $\propto a_{\nu\ell m\lambda}\sigma_{eg}$ channel; only the counterrotating $\propto a^{\dagger}_{\nu\ell m\lambda}\sigma_{eg}$ contributes to excitation. This singles out the terms displayed in (IV.23)–(IV.24) and underlies the emission-before-absorption picture adopted in the master-equation treatment.
@@
    $\nu\geq m_{V},\qquad k_{\infty}(\nu)=\sqrt{\nu^{2}-m_{V}^{2}},$ (IV.25)
    so modes below threshold do not contribute to the escaping flux. In the stationary-phase evaluation, the rapidly varying piece depends only on $\nu>0$, thus preserving the same detailed-balance structure as in Eqs. (B13)-(B15) of [44].
@@
    $\sum_{\lambda}\!\big|u\!\cdot\!\varepsilon^{(\lambda)}\big|^{2}=-1+\frac{\Omega^{2}(\tau,\nu)}{m_{V}^{2}},$ (IV.26)
@@
    Since all master fields reduce to plane waves $\sim e^{-i\nu(t-r_{*})}$ as $r\to r_{g}^{+}$, the same stationary-phase structure that yielded Eq. (B14) in [44] in the scalar case recurs here, while the mass and polarization affect only the prefactor (and the $\nu\geq m_{V}$ threshold).
These rules ensure that, after summing over $(m,\lambda)$ and implementing the single-mode cavity filter, the excitation probabilities reduce to the forms already prepared in (IV.10)–(IV.11) and (IV.21)–(IV.22). When we convert these amplitudes into rates and feed them into the master equation, we will recover the same detailed-balance ratio $\Gamma_{e}/\Gamma_{a}=e^{-2\xi}$ with $\xi=2\pi r_{g}\nu/c$, exactly mirroring the scalar result while carrying the Proca-specific prefactors and the hard mass threshold.
With the emission amplitudes written using $\nu\left[t(\tau)-r_{*}(\tau)\right]$ in the rapidly varying phase, the stationary-phase evaluation repeats Scully et al.’s steps [their (B11)–(B14)] verbatim [44], yielding the same Planckian factor, while the Proca mass/polarization affects only the smooth prefactor and the $\nu\geq m_{V}$ threshold.
@@ V Excitation probability along the infall worldline
We evaluate the first-order excitation amplitudes derived in Sec. IV along the infalling geodesic. For both detector models, the integrals are dominated by the near-horizon segment where the phase is governed by the combination $t-r_{*}$. Using the geodesic relations from Sec. II and the unit-flux mode normalization from Sec. III, the worldline integral reduces, in the large-$\omega$ regime, to an end point (near-horizon) asymptotics that yields a gamma-function. Its modulus square produces the Planckian detailed-balance factor $\left(e^{4\pi\nu}-1\right)^{-1}$. Mass and polarization enter only through smooth prefactors and the asymptotic kinematic threshold $\nu\geq m_{V}$.
@@ V.1 General stationary-phase integral
For a single outgoing mode $(\nu,\ell,m,\lambda)$, the first-order counterrotating amplitude has the structure
$\mathcal{A}_{\bullet}(\nu,\ell,m,\lambda)=\int_{-\infty}^{+\infty}\!d\tau\;\mathcal{G}_{\bullet}(\tau,\nu,\ell,m,\lambda)\,e^{\,i\Phi(\tau,\nu)},$ (V.1)
where $\bullet\in{\text{mono},\text{dip}}$ denotes the detector model. The smooth envelope $\mathcal{G}_{\bullet}$ collects the $1/[r(\tau)\sqrt{k_{\infty}(\nu)}]$ factor from mode normalization and the appropriate polarization weight (cf. (IV.11) for the current coupling and its dipole analogue), while the phase is
$\Phi(\tau,\nu)=\omega\tau+\nu\,\left[t(\tau)-r_{*}(\tau)\right]+\text{(constant)}$ (V.2)
@@
$\frac{dr}{d\tau}=-\frac{1}{\sqrt{r}},\quad\frac{dt}{d\tau}=\frac{r}{r-1},\quad\frac{dr_{*}}{d\tau}=\frac{dr}{d\tau}\,\frac{dr_{*}}{dr}=-\frac{\sqrt{r}}{\,r-1\,}.$ (V.3)
@@
$\frac{d}{d\tau}(t-r_{*})=\frac{r+\sqrt{r}}{r-1}\approx\frac{1+(1+\delta/2)}{\delta}\approx\frac{2}{y}.$ (V.4)
@@
$t(\tau)-r_{*}(\tau)=-\,2\ln y+C_{h}+\mathcal{O}(y),\qquad y\to 0^{+},$ (V.5)
consistent with surface gravity $\kappa=\frac{1}{2}$ in our dimensionless units. Substituting (V.5) into (V.2) and absorbing constants into an overall phase,
$\Phi(\tau,\nu)=\Phi_{h}-\omega y-2\nu\ln y+\Delta(\nu)\,r_{*}(\tau_{h})+\mathcal{O}(y).$ (V.6)
Since $\mathcal{G}_{\bullet}(\tau,\nu,\ell,m,\lambda)$ is smooth, we replace it by its horizon value $\mathcal{G}_{\bullet}(\tau_{h};\nu,\ell,m,\lambda)\equiv\mathcal{G}_{\bullet h}(\nu,\ell,m,\lambda)$ at leading order. The amplitude (V.1) becomes
$\mathcal{A}_{\bullet}(\nu,\ell,m,\lambda)=\mathcal{N}_{\bullet}(\nu,\ell,m,\lambda)\,\underbrace{\int_{0}^{\infty}\!dy\;y^{-2i\nu}\,e^{-i\omega y}}_{\displaystyle\mathcal{I}(\nu,\omega)},$ (V.7)
@@
The elementary Laplace-type integral in (V.7) is expressed in terms of the gamma function,
$I(\nu,\omega)=(i\omega)^{-(1-2i\nu)}\Gamma(1-2i\nu),\qquad\omega>0.$ (V.8)
@@
$|\mathcal{I}(\nu,\omega)|^{2}$ $=\frac{1}{\omega^{2}}\left|\Gamma(1-2i\nu)\right|^{2}=\frac{2\pi\nu}{\omega^{2}}\;\frac{1}{\sinh(2\pi\nu)}$
$=\frac{4\pi\nu}{\omega^{2}}\;\frac{1}{e^{2\pi\nu}-e^{-2\pi\nu}}\;.$ (V.9)
Equation (V.9) governs the thermal character of the problem. While $|\mathcal{I}|^{2}$ itself follows a statistics-dependent profile (here proportional to $\text{cosech}(2\pi\nu)$), the crucial feature for thermodynamics is the relation between excitation (emission) and de-excitation (absorption). All dependence on the detector model (current vs. dipole), the Proca mass and polarizations, and the asymptotic threshold $\nu\geq m_{V}$ resides in the smooth prefactor $\mathcal{N}_{\bullet}$ and in $k_{\infty}(\nu)$. Consequently, at leading order in $1/\omega$, the ratio of probabilities satisfies the detailed balance condition $P_{\text{exc}}/P_{\text{abs}}=e^{-4\pi\nu}$ (derived below via the KMS condition), ensuring the steady state is thermal.
@@ V.2 Universality of the thermal factor
The Planckian denominator obtained in (V.9) is independent of the detector model (current vs. dipole), the field’s spin, the Proca mass, and greybody details. The reason is structural: after isolating a single outgoing Schwarzschild mode, every first-order counterrotating amplitude reduces to an integral of the form
$\int d\tau\,\underbrace{\mathcal{G}(\tau,\nu)}_{\text{smooth}}\,\exp\!\big\{i\omega\tau+i\nu\,\left[t(\tau)-r_{*}(\tau)\right]+i\Delta(\nu)r_{*}(\tau)\Big\},$ (V.10)
@@
$|\mathcal{I}(\nu,\omega)|^{2}=\frac{4\pi\nu}{\omega^{2}}\,\frac{1}{e^{4\pi\nu}-1}\;.$ (V.13)
Equation (V.13) is the universal thermal factor for the excitation probabilities. All dependence on the detector realization (current vs. dipole), polarization content, Proca mass and threshold $\nu\geq m_{V}$, and barrier transmission enters only through the smooth prefactor that multiplies (V.13).
The universality of the excitation mechanism is displayed in Fig. 2, which plots the squared modulus of the near-horizon integral $|\mathcal{I}(\nu)|^{2}$. This profile, governed by the $\text{cosech}(2\pi\nu)$ dependence derived in Eq. (V.9), dictates the thermal response of the detector. Crucially, this curve is identical for both the monopole and dipole couplings and for any mass $m_{V}$, confirming that the thermal character arises solely from the geometric analyticity of the Rindler time coordinate $t-r_{*}$ rather than the specific details of the probe or the field. The detailed balance ratio required for thermal equilibrium emerges from the symmetry of this kernel under $\nu\to-\nu$.
@@
The same near-horizon analyticity that leads to (V.13) also fixes detailed balance between the counterrotating excitation channel and its absorption counterpart (appearing later in the master equation). Under $\nu\to-\nu$ the logarithmic phase in (V.11) implements the (Kubo-Martin-Schwinger) KMS-type periodicity of Rindler time, giving
$P_{\rm abs}(\nu)=e^{\,4\pi\nu}\,P_{\rm exc}(\nu)\;,$ (V.14)
@@ V.3 Prefactors, thresholds, and detailed balance
The near-horizon integral in (V.7)–(V.9) fixes the entire $\omega$- and $\nu$-dependence that carries the thermal factor. Everything else, from detector realization, polarization structure, greybody transmission, and the Proca mass, enters multiplicatively through a smooth prefactor evaluated on the horizon segment, together with the asymptotic kinematic threshold $\nu\geq m_{V}$.
@@
$P_{\text{exc}}^{\bullet}(\nu,\ell)$ $=\frac{4\pi\nu}{\omega^{2}}\;\frac{\mathcal{K}_{\bullet}\,\Xi_{\ell}(\nu)}{e^{4\pi\nu}-1}\;$
$\times\frac{\Theta(\nu-m_{V})}{k_{\infty}(\nu)}\;\overline{\mathcal{P}}_{\bullet}(\nu,\ell)\left[1+\mathcal{O}\!\left(\tfrac{\nu}{\omega}\right)\right],$
$k_{\infty}(\nu)$ $=\sqrt{\nu^{2}-m_{V}^{2}},$ (V.15)
where $\mathcal{K}_{\bullet}$ collects the detector coupling and matrix element $(\mathcal{K}_{\text{mono}}=g_{V}^{2}|m_{eg}|^{2}$, $\mathcal{K}_{\text{dip}}=g_{d}^{2}|d_{eg}|^{2})$, $\Xi_{\ell}(\nu)$ is the single-channel transmission (greybody) through the exterior potential barrier, $k_{\infty}(\nu)^{-1}$ comes from the unit-flux mode normalization at infinity, and $\overline{\mathcal{P}}_{\bullet}(\nu,\ell)$ is a smooth, order-unity polarization/geometry factor fixed by the corresponding contraction in Sec. IV and by the cavity’s angular selection.
Specializing (V.15) to the two detector models requires handling the sum over $m$. For an atom falling radially along the symmetry axis ($\theta\approx 0$), the interaction is weighted by $Y_{\ell m}(\theta\approx 0)$, which vanishes for $m\neq 0$. Consequently, the sum over $m$ collapses to the single $m=0$ term. With this selection, the probabilities become
$P_{\text{exc}}^{\text{(mono)}}(\nu,\ell)$ $=\frac{4\pi\nu}{\omega^{2}}\;\frac{g_{V}^{2}|m_{eg}|^{2}}{e^{4\pi\nu}-1}\;\frac{\Xi_{\ell}(\nu)}{k_{\infty}(\nu)}\;$
$\times\Theta(\nu-m_{V})\;\overline{\mathcal{P}}_{\text{mono}}(\nu,\ell,m=0)\;,$ (V.16)
$P_{\text{exc}}^{\text{(dip)}}(\nu,\ell)$ $=\frac{4\pi\nu}{\omega^{2}}\;\frac{g_{d}^{2}|d_{eg}|^{2}}{e^{4\pi\nu}-1}\;\frac{\Xi_{\ell}(\nu)}{k_{\infty}(\nu)}\;$ (V.17)
$\times\Theta(\nu-m_{V})\;\overline{\mathcal{P}}_{\text{dip}}(\nu,\ell,m=0)\;.$ (V.18)
@@
The hard mass threshold is enforced by $\Theta(\nu-m_{V})$; close to threshold the asymptotic momentum $k_{\infty}(\nu)$ suppresses the probability as $k_{\infty}^{-1}$, while the greybody factor $\Xi_{\ell}(\nu)$ further damps low $\nu$. In the massless limit $m_{V}\to 0$ one has $k_{\infty}\to\nu$ so the prefactor tends to the familiar $1/\nu$ normalization, reproducing the scalarlike spectral scaling at fixed $\Xi_{\ell}(\nu)$ and $\overline{\mathcal{P}}_{\bullet}$.
Detailed balance follows from the same near-horizon analyticity that produced (V.13). Replacing $\nu\to-\nu$ in the phase reverses the sign of the logarithm and yields
$P_{\text{abs}}^{\bullet}(\nu,\ell)=e^{\,4\pi\nu}\,P_{\text{exc}}^{\bullet}(\nu,\ell),\qquad\nu>0,$ (V.19)
@@
$\frac{P_{\text{exc}}^{\bullet}(\nu,\ell)}{P_{\text{abs}}^{\bullet}(\nu,\ell)}=\frac{1}{e^{\,4\pi\nu}}\;.$ (V.20)
Equations (V.16)–(V.20) are the inputs we will promote to rates in the next section by dividing by the interaction time set by the cavity transit and by incorporating the single-mode selection, thereby closing the loop to the master-equation description.